make m the subject of the formula W=(F/M+g/n)^-1
please explain to me how do you do it dont just give me the answer.

Answer by stanbon(75887)

make m the subject of the formula W=(F/M+g/n)^-1
----
Invert both sides to get:
(F/M+g/n) = 1/W
----
Subtract g/n from both sides:
F/M = (1/W)-(g/n)
---
Simplify the right side:
F/M = (n-gW)/(nW)
---
Invert both sides:
M/F = (nW)/(n-gW)
---
Multiply both sides by "F":
M = (FnW)/(n-gW)
======================
Cheers,
Stan H.